Supplemental Operations Report – March 9th

Supplemental Operations Report 9 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine (full week report, I missed last Sunday)

Non-nominal systems: Outpost water system, Sojourner

Action taken for non-nominal systems: I am emptying the water supply as much as possible, then I will shock it with bleach and refill the storage tank. I’m too big to get into the tank to clean it. I don’t know for sure that the water is bad but we have been having stomach issues and it was not cleaned at the beginning of the field season so this is just in case.

Sojourner’s batteries are not working properly and are losing charge much quicker than when I left in January. I will be checking and refilling all battery cells and see if that improves the power.

Generator: Still being manually run at night. Approximate hours 9 pm to 7 am on sunny days, slightly longer (with an earlier start time at night) if it is cloudy.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— SOC is nominal.

Notes on power system: Nothing new to report. I will be working on the system this weekend, attempting to get the generator to charge the batteries at night.

Propane Reading, station tank – 64 %

Propane Reading, director tank— 70+ % (estimated at 80%)

Propane Reading, intern tank— 80 %

Propane Reading, generator— 70 %

Ethanol Free Gasoline – not known gallons.

Water (Outpost tank) – 75 gallons

Sojourner rover used: yes

Hours: 162.4

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Sojourner is losing charge far too quickly.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs:

HabCar used and why, where? Yes, to town for garbage runs and supplies

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

Campus-wide inspection, if action is taken, what and why? Nothing to report

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing new to report.

Summary of suits and radios: One helmet has damage and has been switched for the spare one at the Outpost.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Greywater tank valve needs replacement.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: We tried to open the observatory dome using an onsite internet connection but failed. We will attempt manually opening it next week.

Summary of health and safety issues: Possible contamination of outpost water supply.

Copyright © The Mars Society. All rights reserved. | Main Site